1. Two charges Q1=-5 NC and Q2=-7 nc are separated by 8 cm. Calculate tHe electric potential at point.

2. A research van de Graaf generator has a 2.50 mm diameter sphere with a charge of 6.00 UC on it.

 A. What is the potential near its surface?

 B. At what distance from its center is the potential 1.50 MV?


1
Expert's answer
2021-04-26T17:12:47-0400

1. The electric potential at the middle is


"\phi=\frac{kq_1}{r}+\frac{kq_2}{r}=-2696\text{ V}."

2. The potential near surface is


"\phi_A=\frac{kq}{r}=43.14·10^6\text{ V}."

The potential is 1.5 MV is at the following distance from the center:


"r=\frac{kq}{\phi_B}=0.036\text{ m}."